hartnn (hartnn):

Show that \( \Large (1-e^{iθ} )^{(-1/2)}+(1-e^{(-iθ)} )^{(-1/2)}=(1+cosec(θ/2) )^{(1/2)} \)

hartnn (hartnn):

\(\large (1-e^{iθ} )^{(-1/2)}+(1-e^{(-iθ)} )^{(-1/2)}=(1+cosec(θ/2) )^{(1/2)}\)

hartnn (hartnn):

started with right side, couldn't go too far...now trying to start with left side

zepdrix (zepdrix):

Hmm I can't seem to figure it out D: I also tried starting from the right side. Starting by using this:\[\Large\rm \sin \theta=\frac{e^{i \theta}-e^{-i \theta}}{2i}\]And then after the calculations, turning it into (1+csc(theta/2))^(1/2) gives us,\[\Large\rm \left(1+\csc\frac{\theta}{2}\right)^{1/2}=\left(1+\frac{2i}{(e^{i \theta})^{1/2}-(e^{-i \theta})^{1/2}}\right)^{1/2}\]But I got stuck there >.< Mmmm

hartnn (hartnn):

\( \Large \dfrac{(2i \sin θ)}{((\sqrt{(1-e^{(-iθ)} )}-\sqrt {(1-e^{iθ} ))} \sqrt {(2-e^{(-iθ)}-e^{iθ} )}) } \) me here

OpenStudy (shadowlegendx):

Hartnn someday you need to teach me 1 + 1, you're so intelligent o-o

hartnn (hartnn):

sorry, i can't teach 1+1, i am too dumb for it :P

hartnn (hartnn):

i am more used to 1+1 =10, instead of 1+1 =2 :P

OpenStudy (larseighner):

Are you sure the denominator is not -1?

hartnn (hartnn):

denominator of what where which side ? :O

OpenStudy (larseighner):

Left. You are multiplying by the conjugate right? I get \[ 1-e^{i\theta} - (1 - e^{-i\theta}) \]

OpenStudy (larseighner):

Oh never mind

hartnn (hartnn):

yep, my left side work, till now is correct, but stuck!

hartnn (hartnn):

\( (2−e^{(−iθ)}−e^{iθ}) = 2-2 cos θ\)

hartnn (hartnn):

any other approach... ?

OpenStudy (thomas5267):

I haven't learn complex number but could you square both sides together?

hartnn (hartnn):

i need to prove left side = right side, so i can't already assume it to be true...

hartnn (hartnn):

anyways, what good will squaring both sides do ?

OpenStudy (thomas5267):

\(\text{LHS}^2=\text{RHS}^2\) is a necessary but not sufficient condition for \(\text{LHS}=\text{RHS}\).

hartnn (hartnn):

the only application of complex numbers here is \(e^{ix } = cos x +i \sin x~,~~ so~~e^{-ix } = cos x -i \sin x \)

OpenStudy (thomas5267):

So theta is real?

hartnn (hartnn):

yep, its an angle...like 30 degree, 60, 90, ...

OpenStudy (thomas5267):

Am I allowed to do binomial expansion on the left hand side if I take the square of both sides?

hartnn (hartnn):

yes, you can use any math formula available

OpenStudy (thomas5267):

Something is wrong in here. \[ \begin{align*} &\left(\left(1-e^{i\theta}\right)^{-1/2}+\left(1-e^{-i\theta}\right)^{-1/2}\right)^2\\ =&\left(1-e^{i\theta}\right)^{-1}+\left(1-e^{-i\theta}\right)^{-1}+2\left(1-e^{i\theta}\right)^{-1/2}\left(1-e^{-i\theta}\right)^{-1/2}\\ =&\frac{1}{1-e^{i\theta}}+\frac{1}{1-e^{-i\theta}}+2 \\&\text{I couldn't think of a situation that }2\neq2\left(1-e^{i\theta}\right)^{-1/2}\left(1-e^{-i\theta}\right)^{-1/2}\\ =&\frac{1-e^{-i\theta}}{\left(1-e^{i\theta}\right)\left(1-e^{-i\theta}\right)}+\frac{1-e^{i\theta}}{\left(1-e^{i\theta}\right)\left(1-e^{-i\theta}\right)}+2\\ =&\frac{2-e^{i\theta}-e^{-i\theta}}{1-e^{i\theta}-e^{-i\theta}+1}+2\\ =&1+2\qquad???? \end{align*} \]

OpenStudy (anonymous):

From @zepdrix's work... I'm wondering if this gets you somewhere. \[\begin{align*}1+\csc\frac{\theta}{2}&=1+\frac{2i}{\exp\left(\dfrac{i\theta}{2}\right)-\exp\left(-\dfrac{i\theta}{2}\right)}\\\\&=\frac{\exp\left(\dfrac{i\theta}{2}\right)+2i-\exp\left(-\dfrac{i\theta}{2}\right)}{\exp\left(\dfrac{i\theta}{2}\right)-\exp\left(-\dfrac{i\theta}{2}\right)}\\\\ &=\frac{\left(\exp\left(\dfrac{i\theta}{4}\right)+i\exp\left(-\dfrac{i\theta}{4}\right)\right)^2}{\exp\left(\dfrac{i\theta}{2}\right)-\exp\left(-\dfrac{i\theta}{2}\right)}\\\\ \sqrt{1+\csc\frac{\theta}{2}}&=\frac{\exp\left(\dfrac{i\theta}{4}\right)+i\exp\left(-\dfrac{i\theta}{4}\right)}{\sqrt{\exp\left(\dfrac{i\theta}{2}\right)-\exp\left(-\dfrac{i\theta}{2}\right)}}\\\\ \end{align*}\]

hartnn (hartnn):

i got there too, but stuck

hartnn (hartnn):

thomas, 1 is correct, but \(2\neq2\left(1-e^{i\theta}\right)^{-1/2}\left(1-e^{-i\theta}\right)^{-1/2}\\\)

hartnn (hartnn):

ok, \(\sqrt {1+2 \sqrt{2-2\cos \theta }}\) this might lead to right side..

hartnn (hartnn):

\(\sqrt {1+2/[\sqrt 2 \sqrt {(2\sin^2 (\theta/2 ))}]} = \sqrt {1+1/\sin (\theta /2)} =\sqrt {1+cosec \theta/2}\) thanks thomas! :D

OpenStudy (thomas5267):

Where did this \(\sqrt {1+2 \sqrt{2-2\cos \theta }}\) come from?

hartnn (hartnn):

i meant \(\sqrt {1+2/ \sqrt{2-2\cos \theta }}\)

hartnn (hartnn):

because \((1-e^{i\theta })(1-e^{-i\theta } ) = (2-e^{i\theta }-e^{-i\theta })=2-2\cos \theta \)

hartnn (hartnn):

\(\Large\rm \sin \theta=\frac{e^{i \theta}-e^{-i \theta}}{2i} \\ \Large\rm \cos \theta=\frac{e^{i \theta}+e^{-i \theta}}{2}\)

hartnn (hartnn):

you get those ^^ by adding and subtracting these \(e^{ix } = cos x +i \sin x~,~~ so~~e^{-ix } = cos x -i \sin x\)

OpenStudy (thomas5267):

Finally understand what is happening. Is \(\sqrt{a}\sqrt{b}=\sqrt{ab}\text{,}\:a,b\in\mathbb{C}\) generally true?

OpenStudy (larseighner):

Nothing new here. Just thinking aloud. \[\large \begin{align} (1-e^{i\theta} )^{-{1 \over 2}}+(1-e^{-i\theta} )^{-{1 \over 2}}&=(1+\csc(\theta/2) ^{1 \over 2} \cr {1 \over {\sqrt{1-e^{i\theta}}}}+ {1 \over {\sqrt{1-e^{-i\theta}}}} &= \cr { { \sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}} } \over {\sqrt{1-e^{i\theta}}\sqrt{1-e^{-i\theta}}}} &= \cr { { \sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}} } \over { \sqrt{ (1-e^{i\theta})(1-e^{-i\theta} ) } }} &= \cr { { \sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}} } \over { \sqrt{ (1-\cos\theta-i\sin\theta)(1-\cos\theta + i\sin\theta ) } }} &= \cr { { \sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}} } \over { \sqrt{ (2-2\cos\theta ) } }} &= \end{align}\]

OpenStudy (larseighner):

\[ \large \begin{align} { { \sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}} } \over { \sqrt{ (2-2\cos\theta ) } }} &= \cr \left[ { { \sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}} } \over { \sqrt{ (2-2\cos\theta ) } }} \right] ^2 &= 1+\csc{\theta \over 2} \cr { { (\sqrt{1-e^{-i\theta}} +\sqrt{1-e^{i\theta}})^2 } \over { 2-2\cos\theta }} &= \cr { { 1-e^{-i\theta} +2\sqrt{1-e^{-i\theta}}\sqrt{1-e^{i\theta}} + 1-e^{i\theta} } \over { 2-2\cos\theta }} &= \end{align}\]

OpenStudy (larseighner):

\[\large \begin{align} { { 2-2\cos\theta +2\sqrt{1-e^{-i\theta}}\sqrt{1-e^{i\theta}} } \over { 2-2\cos\theta }} &= \cr {{2-2\cos\theta } \over { 2-2\cos\theta }}+{ { 2\sqrt{1-e^{-i\theta}}\sqrt{1-e^{i\theta}} } \over { 2-2\cos\theta }} &= \cr 1 +{ { 2\sqrt{1-e^{-i\theta}}\sqrt{1-e^{i\theta}} } \over { 2-2\cos\theta }} &= 1+\csc{\theta \over 2}\cr { { 2\sqrt{1-e^{-i\theta}}\sqrt{1-e^{i\theta}} } \over { 2-2\cos\theta }} &= \csc{\theta \over 2}\end{align}\]

hartnn (hartnn):

you still on it ? want to see how i solved it ?

OpenStudy (larseighner):

Sure

OpenStudy (larseighner):

clearly the way to go.

OpenStudy (phi):

write \[ \exp(i \theta) = \cos(\theta) + i \sin(\theta) \\ \exp(-i \theta) = \cos(\theta) - i \sin(\theta)\] and \[ 1-\exp(i \theta) = 1 -\cos(\theta) - i \sin(\theta) \\ 1-\exp(-i \theta) = 1 -\cos(\theta) + i \sin(\theta) \] let \[ (1- 1 -\cos(\theta)= a \\ \sin(\theta)= b \] the left-hand side is \[ \frac{1}{\sqrt{a- bi}} + \frac{1}{\sqrt{a+bi}} \\ = \frac{\sqrt{a+bi}+\sqrt{a-bi}}{\sqrt{a^2+b^2}} \] we use \[ a^2 + b^2 = 1+\cos^2(\theta)- 2 \cos(\theta)+\sin^2(\theta) = 2(1- \cos(\theta)= 2a\] to write \[ \frac{\sqrt{a+bi}+\sqrt{a-bi}}{\sqrt{2a}} \] square this quantity \[ \frac{a+bi + a - bi + 2\sqrt{a^2+b^2}}{2a} \\ =\frac{ 2a+2 \sqrt{2a} }{2a} \\ = 1 + \frac{2}{\sqrt{2a}} \] meanwhile, the square of the right -hand side is \[ 1+ \frac{1}{\sin(\frac{x}{2})} \\ = 1 + \frac{1}{\sqrt{\frac{1}{2} (1-\cos(x))}}\\ = 1 + \frac{1}{\sqrt{\frac{1}{2} a}}\\ = 1 + \frac{2}{\sqrt{2a}} \] which matches the left-hand side

hartnn (hartnn):

yep, no other way squaring sides is necessary to prove this!

OpenStudy (phi):

**typo let \[ 1- \cos(\theta)= a \\ \sin(\theta)= b \]

OpenStudy (phi):

yes, it took some putzing around to find a way out of the weeds.